OpenStudy (anonymous):

S = p q + p(1 − q) for p

OpenStudy (anonymous):

\[S = pq + p(1-q)\]Distribute out the p(1-q): \[S = pq + p - pq\]Combine like-terms: \[S = p\]

OpenStudy (anonymous):

Thanks!